# Turbulence, non-equilibrium thermodynamics

### v1

@@ 3 Fluctuation relation analysis from single-time data
with a slope of order unity (about 1.13 was found in Yao et al. (2023)).
In the present study, we utilize direct numerical simulations of forced isotropic turbulence from the publicly available Johns Hopkins Turbulence Database (JHTDB; Yao et al. (2025)). As in Yao et al. (2023), the analysis is conducted at three length scales within the inertial range: $\ell=30\eta$, $45\eta$, and $60\eta$, where $\eta$ denotes the Kolmogorov length scale computed based on the overall rate of dissipation of the flow. To compute the surface averages required for evaluating $\Phi_{\ell}$, the outer surface of diameter $\ell$ is discretized into 500 approximately uniformly distributed point pairs (denoted as $+$ and -- points) on a sphere. The velocity components used to calculate $\delta u_{i}$ are obtained via the JHTDB function ‘getData’ using an 8th-order Lagrangian spatial interpolation scheme. The quantity $k_{\ell}$ is similarly computed by a discretized volume integration by integrating over five concentric spherical shells. The accuracy of this surface and volume integration approach has been verified through convergence tests by increasing the number of discretization points (Yao et al., 2024a). The quantities $k_{\ell}$, $\Phi_{\ell}$, and the rate $\Phi_{\ell}/k_{\ell}$ (proportional to the entropy generation rate) are computed at one million spheres centered on random points, sampled across 2000 spatial locations within the three-dimensional domain $[0,2\pi]^{3}$, over 500 statistically stationary time snapshots during the available 5 large-scale turn-over time interval available in the database. Although $\Phi_{\ell}/k_{\ell}$ is computed across multiple time snapshots for statistical convergence of the PDFs, the present method does not involve any temporal averaging. Given the statistical stationarity in time and homogeneity in space characteristic of isotropic turbulence, each sample can be regarded as an independent realization of the local, instantaneous flow state. As a result, the combined set of spatio-temporal samples constitutes a valid statistical ensemble for constructing the PDF of $\Phi_{\ell}/k_{\ell}$ and and for testing the constant-time FR.
Results are shown in Figure 1. A close to linear relationship is observed for the logarithm of the probability density ratio of positive and negative entropy generation rate, at the same rate across the three length scales. The dashed line has a slope of 1.16, very close to the earlier result at higher Reynolds number, and confirming that the FR as observed in Yao et al. (2023) is valid also at this lower Reynolds number.
@@
4 Finite time-averaged Fluctuation Relation
@@ 4.1 Lagrangian framework
Next, we examine the PDFs of the finite time-averaged $\Psi^{\tau}_{\ell}$. In previous work (Yao et al., 2023, 2024b), the PDF of the entropy generation rate at a single time was found to be very closely exponential, on both the positive and negative sides. For such two-sided exponential PDFs, the slope of the linear relationship (in the semi-log plots) demonstrating the FR is simply the difference between the negative and positive slopes (Yao et al., 2023). In contrast, for the finite-time-averaged case, we observe a qualitatively different behavior, as shown in Figure 5. For positive values the PDF is curved and close to Gaussian (non-exponential) while at negative values, as the averaging time $\tau$ increases, the PDF gradually shifts from an exponential shape to a more parabolic (or Gaussian-like) profile. The PDFs have a long, non-Gaussian tails but near the peak, they are more curved, which means that the low-intensity values behave more Gaussian-like than the extreme values at the tails of the PDFs. This transition is consistent with what was reported in several original FR studies for sheared systems, where the PDF of the entropy-like variable becomes Gaussian in the long-time limit (figure 1 of Evans et al. (1993)). One can attribute this type of trend to the central limit theorem, where increasing $\tau$ effectively aggregates more independent events (e.g., intervals of $\tau$ with non-Gaussian PDFs), leading towards Gaussian behavior when $\tau$ becomes large. Interestingly, regardless of how the PDFs evolve in shape (from one-sided exponential to Gaussian), the FR remains valid.
@@ 4.2 Eulerian framework
The failure of the FR when analyzing data in an Eulerian perspective confirms the importance of considering spatial advection of turbulent fluid parcels. Since the definition of entropy of turbulence is expected to be associated with a material region moving with the flow at scale $\ell$, the observation that the FR holds only when properly following the flow provides support for the definition of entropy of turbulence used in this and prior works (Yao et al., 2023).
@@ 5 Conclusion
In this work, we have extended the concept of entropy generation rate in turbulent flows by redefining the original definition proposed in Yao et al. (2023). The key modification involves the definition of the temperature of turbulence, taken to be proportional to the turbulent kinetic energy. Unlike the kinetic theory of gases, where the number of degrees of freedom associated with each microscopic particle is known, in the case of turbulence we do not know this value. Hence it was now left as an empirical parameter $\gamma_{t}$.
@@
We first tested the constant-time approximation to the FR using data from forced isotropic turbulence at a moderate Reynolds number, $Re_{\lambda}=433$ from JHTDB. We confirmed that the linear trend predicted from the FR holds at this lower-Reynolds number dataset yielding essentially the same results as the higher Reynolds number data considered in Yao et al. (2023). Next, we examined the more fundamental finite time-averaged FR under a Lagrangian framework, where time averaging is performed along trajectories defined by the spatially filtered velocity field. Our results demonstrate that the FR holds robustly, i.e. the results show linear behavior in semi-logarithmic plots of probability ratios versus entropy generation. Various averaging time-intervals were tested, ranging up to $3\tau_{\ell}$. In contrast, when the same analysis was performed in an Eulerian framework, the FR no longer holds and the ratio of probabilities no longer follows exponential behavior. This breakdown of the FR might be attributed to the lack of trajectory/temporal coherence, which violates the conditions required for the definition of entropy as relating to fluid material.
@@
While we have demonstrated that the FR holds under filtered Lagrangian trajectories, it remains an open question whether this behavior is unique to the specific choice of trajectory. Future studies should explore alternative definitions, such as trajectories based on unfiltered velocities, filtered velocities in different directions, or flow-aligned coherent structures. Additionally, the connection between physical space trajectories and phase-space trajectories, as described in Fokker–Planck or Liouville frameworks (Fuchs et al., 2020), warrants deeper exploration, particularly in the context of finite-time Lyapunov exponents and attractor dynamics. Finally, it would be natural to extend this study to more complex turbulent systems, including anisotropic turbulence, shear flows, wall-bounded turbulence, and two-dimensional turbulence, where different mechanisms of entropy production and time asymmetry are expected to arise. It will also be of interest to repeat the analysis using Large-Eddy-Simulation (LES) versions of the energy flux and temperature, for which the single-time analysis of Yao et al. (2023) showed strong deviations from the FR.
[…]
@@
We are grateful to the JHTDB/IDIES staff for their assistance with the database and its maintenance. This work is supported by NSF (Grant # CSSI-2103874).

### v2

@@ III Fluctuation relation analysis from single-time data
with a slope of order unity (about 1.13 was found in [1]).
In the present study, we utilize direct numerical simulations of forced isotropic turbulence from the publicly available Johns Hopkins Turbulence Database (JHTDB; [30]). As in [1], the analysis is conducted at three length scales within the inertial range: $\ell=30\eta$, $45\eta$, and $60\eta$, where $\eta$ denotes the Kolmogorov length scale computed based on the overall rate of dissipation of the flow. To compute the surface averages required for evaluating $\Phi_{\ell}$, the outer surface of diameter $\ell$ is discretized into 500 approximately uniformly distributed point pairs (denoted as $+$ and -- points) on a sphere. The velocity components used to calculate $\delta u_{i}$ are obtained via the JHTDB function ‘getData’ using an 8th-order Lagrangian spatial interpolation scheme. The quantity $k_{\ell}$ is similarly computed by a discretized volume integration by integrating over five concentric spherical shells. The accuracy of this surface and volume integration approach has been verified through convergence tests by increasing the number of discretization points [29]. The quantities $k_{\ell}$, $\Phi_{\ell}$, and the rate $\Phi_{\ell}/k_{\ell}$ (proportional to the entropy generation rate) are computed at one million spheres centered on random points, sampled across 2000 spatial locations within the three-dimensional domain $[0,2\pi]^{3}$, over 500 statistically stationary time snapshots during the available 5 large-scale turn-over time interval available in the database. Although $\Phi_{\ell}/k_{\ell}$ is computed across multiple time snapshots for statistical convergence of the PDFs, the present method does not involve any temporal averaging. Given the statistical stationarity in time and homogeneity in space characteristic of isotropic turbulence, each sample can be regarded as an independent realization of the local, instantaneous flow state. As a result, the combined set of spatio-temporal samples constitutes a valid statistical ensemble for constructing the PDF of $\Phi_{\ell}/k_{\ell}$ and and for testing the constant-time FR.
Results are shown in Figure 1. A close to linear relationship is observed for the logarithm of the probability density ratio of positive and negative entropy generation rate, at the same rate across the three length scales. The dashed line has a slope of 1.16, very close to the earlier result at higher Reynolds number, and confirming that the FR as observed in [1] is valid also at this lower Reynolds number.
@@
IV Finite time-averaged Fluctuation Relation
@@ IV.1 Lagrangian framework
Next, we examine the PDFs of the finite time-averaged $\Psi^{\tau}_{\ell}$. In previous work [1, 26], the PDF of the entropy generation rate at a single time was found to be very closely exponential, on both the positive and negative sides. For such two-sided exponential PDFs, the slope of the linear relationship (in the semi-log plots) demonstrating the FR is simply the difference between the negative and positive slopes [1]. In contrast, for the finite-time-averaged case, we observe a qualitatively different behavior, as shown in Figure 5. For positive values the PDF is curved and close to Gaussian (non-exponential) while at negative values, as the averaging time $\tau$ increases, the PDF gradually shifts from an exponential shape to a more parabolic (or Gaussian-like) profile. The PDFs have a long, non-Gaussian tails but near the peak, they are more curved, which means that the low-intensity values behave more Gaussian-like than the extreme values at the tails of the PDFs. This transition is consistent with what was reported in several original FR studies for sheared systems, where the PDF of the entropy-like variable becomes Gaussian in the long-time limit (figure 1 of [3]). One can attribute this type of trend to the central limit theorem, where increasing $\tau$ effectively aggregates more independent events (e.g., intervals of $\tau$ with non-Gaussian PDFs), leading towards Gaussian behavior when $\tau$ becomes large. Interestingly, regardless of how the PDFs evolve in shape (from one-sided exponential to Gaussian), the FR remains valid.
@@ IV.2 Eulerian framework
The failure of the FR when analyzing data in an Eulerian perspective confirms the importance of considering spatial advection of turbulent fluid parcels. Since the definition of entropy of turbulence is expected to be associated with a material region moving with the flow at scale $\ell$, the observation that the FR holds only when properly following the flow provides support for the definition of entropy of turbulence used in this and prior works [1].
@@ V Conclusion
In this work, we have extended the concept of entropy generation rate in turbulent flows by redefining the original definition proposed in [1]. The key modification involves the definition of the temperature of turbulence, taken to be proportional to the turbulent kinetic energy. Unlike the kinetic theory of gases, where the number of degrees of freedom associated with each microscopic particle is known, in the case of turbulence we do not know this value. Hence it was now left as an empirical parameter $\gamma_{t}$.
@@
We first tested the constant-time approximation to the FR using data from forced isotropic turbulence at a moderate Reynolds number, $Re_{\lambda}=433$ from JHTDB. We confirmed that the linear trend predicted from the FR holds at this lower-Reynolds number dataset yielding essentially the same results as the higher Reynolds number data considered in [1]. Next, we examined the more fundamental finite time-averaged FR under a Lagrangian framework, where time averaging is performed along trajectories defined by the spatially filtered velocity field. Our results demonstrate that the FR holds robustly, i.e. the results show linear behavior in semi-logarithmic plots of probability ratios versus entropy generation. Various averaging time-intervals were tested, ranging up to $3\tau_{\ell}$. In contrast, when the same analysis was performed in an Eulerian framework, the FR no longer holds and the ratio of probabilities no longer follows exponential behavior. This breakdown of the FR might be attributed to the lack of trajectory/temporal coherence, which violates the conditions required for the definition of entropy as relating to fluid material.
@@
While we have demonstrated that the FR holds under filtered Lagrangian trajectories, it remains an open question whether this behavior is unique to the specific choice of trajectory. Future studies should explore alternative definitions, such as trajectories based on unfiltered velocities, filtered velocities in different directions, or flow-aligned coherent structures. Additionally, the connection between physical space trajectories and phase-space trajectories, as described in Fokker–Planck or Liouville frameworks [15], warrants deeper exploration, particularly in the context of finite-time Lyapunov exponents and attractor dynamics. Finally, it would be natural to extend this study to more complex turbulent systems, including anisotropic turbulence, shear flows, wall-bounded turbulence, and two-dimensional turbulence, where different mechanisms of entropy production and time asymmetry are expected to arise. It will also be of interest to repeat the analysis using Large-Eddy-Simulation (LES) versions of the energy flux and temperature, for which the single-time analysis of [1] showed strong deviations from the FR.
Acknowledgements.
@@
WWe are grateful to the JHTDB/IDIES staff for their assistance with the database and its maintenance. This work is supported by NSF (Grant # CSSI-2103874).